A body is projected with a velocity 6i+8j+20k i and j being in horizontal plane and k in vertical plane. Horizontal range of the body is a.80 b.34 c.40

Grade:12

1 Answers

Adarsh
6 years ago
Here time of flight = 2.20/10= 4sNet horizontal velocity = √ (8)^2 +(6)^2 = 10m/s Now, net horizontal displacement = 10 . 4 = 40 m Hence (c) is the correct option